OpenStudy (anonymous):

1.04x + 3.14 = 7.98 A. x ≈ 4.6 B. x ≈ 4.5 C. x ≈ 10.7 D. x ≈ 4.7

OpenStudy (anonymous):

is it b

OpenStudy (mathstudent55):

No. Let's do it step by step.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

you have to round to the tenths place

OpenStudy (anonymous):

thats why i said b

OpenStudy (mathstudent55):

1.04x + 3.14 = 7.98 The first step is to subtract 3.14 from both sides. 1.04x + 3.14 = 7.98 - 3.14 -3.14 ------------------- 1.04x = 4.84 Ok so far?

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

Now we divide both sides by 1.04 1.04x/1.04 = 4.84/1.04 x = 4.6538

OpenStudy (calculusxy):